### v3

# Modules of finite injective dimension

Abstract.
Motivated by Bass’ conjecture, we study finitely generated modules of finite injective dimension and the additional constraints they impose on the ambient ring. Beyond ensuring the Cohen–Macaulay property, the presence of such modules enforces further conditions on the ring, including reducedness, normality, being domain, and various (non)-singularity conditions such as complete intersection, Gorenstein, and beyond. We also address the reflexivity, (and also torsionlessness), of modules with finite injective dimension and show that this forces the ring to be quasi-normal.
Key words and phrases:
Cohen-Macualay rings; injective dimension; integral domains; regular (Gorenstein) rings; Auslander’s zero-divisor; regular sequences.
2020 Mathematics Subject Classification:
13C14; 13H10; 13D22
1. Introduction
Bass [10] conjectured that the existence of a finitely generated module of finite injective dimension imposes strong restrictions on a ring, specifically requiring it to be Cohen–Macaulay. Peskine and Szpiro [32] verified the conjecture in many cases, and it was ultimately resolved in full generality by Roberts [33]. Despite its resolution, the statement remains commonly referred to as Bass’ conjecture. The work of Peskine and Szpiro also includes numerous related results of independent interest. In particular, they investigated additional constraints on the ring when attention is restricted to certain classes of modules with finite injective dimension. We also note that Matlis studied ideals of injective dimension one. These considerations motivate the present note.
Let $\mathfrak{p}$ be in $\operatorname{Spec}(R)$, and suppose $\operatorname{pd}_{R}(\mathfrak{p})<\infty$, equivalently $\operatorname{pd}_{R}(R/\mathfrak{p})<\infty$. Peskine-Szpiro [32, Corollary II.3.3] proved that $R$ is an integral domain. It is easy to see that the ring is a domain if $\operatorname{id}_{R}(R/\mathfrak{p})<\infty$. This leads to the following question:
Question 1.1.
If $R$ admits a prime ideal $\mathfrak{p}$ such that $\operatorname{id}_{R}(\mathfrak{p})<\infty$, what consequences does this have for the structure of $R$?
Some of the results from Section 2 are summarized in the following items:
Observation 1.2.
  1. 1)
    Suppose that $\operatorname{id}_{R}(\mathfrak{p})<\infty$ for all $\mathfrak{p}\in\operatorname{Assh}(R)$. Then $R$ is reduced.
  2. 2)
    Suppose $\min(R)$ is singleton. If $\operatorname{id}_{R}(\mathfrak{p})<\infty$ for some $\mathfrak{p}\in\operatorname{Spec}(R)$, then $R$ is a domain.
  3. 3)
    Suppose that $\operatorname{id}_{R}(\mathfrak{p})<\infty$ for all $\mathfrak{p}\in\operatorname{Spec}(R)$ of height at most one. Then $R$ is normal.
  4. 4)
    Let $R$ be an excellent $\operatorname{UFD}$ containing $k$ and of dimension greater than 3. Suppose that $\operatorname{id}_{R}(\mathfrak{p})<\infty$ for all $\mathfrak{p}\in\operatorname{Spec}(R)$ of height two. Then $R$ is regular.
We also present a quasi-version of the reduced and normal items. For example, we show:
Corollary 1.3.
Suppose that $\operatorname{id}_{R}(\mathfrak{p})<\infty$ for some $\mathfrak{p}\in\operatorname{Spec}(R)$ of positive height. Then $R$ is quasi-normal.
Suppose we allow more ideals to have finite injective dimension. For instance, suppose that $\operatorname{id}_{R}(\mathfrak{a})<\infty$ for all ideals of big-height at most two. Then, applying our displayed results plus a famous result of Auslander [6], we show that $R$ is regular.
In Section 3, we focus on modules of finite injective dimension with few generators. In fact, Vasconcelos asked in [36, Page 51], how do Gorenstein rings arise? and studied the question by looking at certain modules with few generators, see [36, 2.40]. Our first result in §3 is as follows:
Proposition 1.4.
Let $(R,\mathfrak{m})$ be quasi-normal, and let $M$ be such that:
  1. i)
    $\operatorname{id}_{R}(M)<\infty$,
  2. ii)
    $M$ is reflexive,
  3. iii)
    $\mu(M)\leq 2$.
Then $R$ is Gorenstein.
As an application, we extend Proposition 1.4 to the case of four-generated modules over rings of type at most two. We also present some comments on §2. Then, we slightly extend [36, 2.40] to a more general setting.
Suppose $M$ is a nonzero tor-rigid module. Auslander [8, 4.3] proved that each $M$-regular sequence is an $R$-regular sequence. As modules of finite projective dimension are not generally tor-rigid, he conjectured (proved by Roberts [34, 6.2.3]) that the above property holds for modules of finite projective dimension. This is referred to as Auslander’s zero-divisor conjecture. In Section 4, we present the injective analogue of Auslander’s zero-divisor conjecture. We do this via two different methods, which we use to reconstruct Auslander’s zero-divisor conjecture over Cohen-Macaulay rings. This also enables us to reprove some results from §2.
In Section 5, we address the following question (see [20]):
Question 1.5.
Suppose $\operatorname{id}_{R}(I/I^{2})$ is finite. Is $I$ generated by a regular sequence?
We present two cases where the answer is positive, namely in the almost complete-intersection case. This immediately recovers the main result of [25] by Kunz.
Recall that [36, 2.38] suggests that the canonical module is an important source for producing, in a constructive way, modules of finite injective dimension. In fact, Matlis applied this to study the following questions:
Question 1.6.
Let $R$ be a $1$-dimensional integral domain with fraction field $Q$. When is $Q/R\twoheadrightarrow E_{R}(k)\to 0$? When is $Q/I=E_{R}(k)$?
To see how these are related to modules of finite injective dimension, recall that (ii) is equivalent to ideals of finite injective dimension. In Section 6, we extend these results to the higher-dimensional situation. It may be nice to recall that $R$ is called quasi-Gorenstein if $H^{\dim(R)}_{\mathfrak{m}}(R)\cong E_{R}(k)$, so Matlis was also interested in this and its semi-form as well. Then, we extend an observation of Auslander [5] about the cohomology of $Q/R$ and present its converse part, see Observation 6.6.
Section 7 addresses the reflexivity of modules with finite injective dimension. We begin by showing that if a nonzero torsionless module has finite injective dimension, then the ring must be generically Gorenstein. We then extend this result to reflexive modules as follows:
Proposition 1.7.
Suppose there exists a nonzero reflexive module $L$ of finite injective dimension. Then $\widehat{R}$ is quasi-normal.
As an immediate application, we drop the quasi-normality assumption from Proposition 1.4 and its consequences. It is easy and well-known that the existence of a nonzero totally reflexive module of finite injective dimension guarantees that the ring is Gorenstein. By $(G_{i})$ we mean $R_{\mathfrak{p}}$ is Gorenstein for all $\mathfrak{p}\in\operatorname{Spec}(R)$ of height at most $i$. As another application, we show:
Observation 1.8.
Let $R$ be a homomorphic image of a Gorenstein ring. Suppose there is a nonzero $(k+2)$-torsionless module with finite injective dimension. Then $R$ is $(G_{k})$.
Section 8 deals with finiteness of finite injective modules of the forms $\operatorname{Hom}_{R}(M,M)$. This was asked in the papers:
Question 1.9.
(See [18, Question 2.18] and [17, Question 4.9]).
  1. i)
    If there exists a nonzero $R$-module $M$ such that $\operatorname{Hom}_{R}(M,M)$ has finite injective dimension, then is $R$ Gorenstein?
  2. ii)
    Let $M$ be a nonzero RR–module such that the injective dimensions of M and $\operatorname{Ext}^{i}_{R}(M,M)$ are finite for all $0\leq i\leq\dim(R)-0pt(M)$. Is $R$ Gorenstein?
We present several cases where the question has an affirmative answer, for example in the ideal case and torsion-free modules when $\mathbb{Q}\subseteq R$. Also, we present a tiny connection from Question 1.9 to the land of prime characteristic, and we show
Observation 1.10.
Let $R$ be $F-\emph{finite}$ and $\operatorname{F}-\emph{injective}$ with quotient singularity. Then $R$ is regular, provided $\operatorname{id}_{R}(\operatorname{Hom}_{R}({{}^{\operatorname{F}_{n}}\!}R,{{}^{\operatorname{F}_{n}}\!}R))<\infty$.
Section 9 is about:
Question 1.11.
(See [14, Question 4.2]). Suppose $R$ is $d$-dimensional local Cohen-Macaulay of minimal multiplicity with canonical module. Let $M$ be a maximal Cohen-Macaulay module so that ${M}\otimes{M}^{\dagger}\cong\omega_{{R}}$. Then is $M\cong R$ or $M\cong\omega_{{R}}$?
We confirm this for totally reflexive modules, even without any use of minimal multiplicity.
2. A family of ideals of small height
In this note $(R,\mathfrak{m},k)$ is a commutative noetherian local ring, and modules are finitely generated, otherwise specialized. The notation $\operatorname{pd}_{R}(-)$ (resp. $\operatorname{id}_{R}(-)$) stands for the projective (resp. injective) dimension of ()(-).
Notation 2.1.
We denote the set of all associated prime ideals by $\operatorname{Ass}_{R}(-)$.
Proposition 2.2.
Suppose $\min(R)$ is singleton. If for some $\mathfrak{p}\in\operatorname{Spec}(R)$ one has $\operatorname{id}_{R}(\mathfrak{p})<\infty$, then $R$ is an integral domain.
Proof.
Following Bass’ conjecture (see [11, Section 9.6]) $R$ is Cohen-Macaulay. In particular, $\operatorname{Ass}(R)=\min(R).$ Suppose $\min(R)=\{P\}$. By definition, there is an $x\in R$ such that $(0:_{R}x)=P$. Suppose $x\in\mathfrak{m}$. Also, recall that
$x\in\operatorname{Zd}(R)=\bigcup_{\mathfrak{q}\in\operatorname{Ass}(R)}\mathfrak{q}=\bigcup_{\mathfrak{q}\in\min(R)}\mathfrak{q}=P=(0:x).$
In other words, $x^{2}=0$. Since $\min(R)$ is singleton, $P\subseteq\mathfrak{p}$. Since $\operatorname{id}_{R}(\mathfrak{p})<\infty$, and due to [11, Proposition 3.1.9], we know that $\operatorname{id}_{R_{\mathfrak{p}}}(\mathfrak{p}R_{\mathfrak{p}})<\infty$. According to Auslander-Buchsbaum-Serre, $R_{\mathfrak{p}}$ is regular (see [11, Theorem 2.2.7]). Since $R_{\mathfrak{p}}$ is regular its localizations are as well. In particular, $R_{P}=(R_{\mathfrak{p}})_{PR_{\mathfrak{p}}}$ is regular. Recall that $x\in P$. Since $x/1\in PR_{P}$, it should be zero. There is an $y\in R\setminus P$ such that $xy=0$. Since $xy=0$ we have $y\in(0:x)=P$. This contradiction says that $x\notin\mathfrak{m}$. Since the ring is local, $x$ is unit. From this $P=(0:x)=0$. Consequently, $R$ is an integral domain. ∎
We denote the set of all associated prime ideal $\mathfrak{p}$ with the property that $\dim R/\mathfrak{p}=\dim R$, by $\operatorname{Assh}_{R}(-)$.
Proposition 2.3.
Suppose that $\operatorname{id}_{R}(\mathfrak{p})<\infty$ for all ${\mathfrak{p}}\in\operatorname{Assh}(R)$. Then $R$ is reduced.
Proof.
Recall that $R$ is Cohen-Macaulay. In particular, $R$ is equidimensional. Let
$\operatorname{Assh}(R)=\operatorname{Ass}(R)=\min(R)=\{\mathfrak{p}_{1},\ldots,\mathfrak{p}_{t}\}.$
In view of Proposition 2.2 we may assume $t>1$. Let $P\in\min(\bigcap_{i=1}^{t}\mathfrak{p}_{i})$. There is an $i$ such that $\mathfrak{p}_{i}\subseteq P$. It follows that $P$ is minimal over $\mathfrak{p}_{i}$. Thus, $P=\mathfrak{p}_{i}$. Since $\operatorname{id}_{R}(\mathfrak{p}_{i})<\infty$, $\operatorname{id}_{R_{\mathfrak{p}_{i}}}(\mathfrak{p}_{i}R_{\mathfrak{p}_{i}})<\infty$. According to Auslander-Buchsbaum-Serre, $R_{\mathfrak{p}_{i}}$ is regular. Then $(\bigcap_{i=1}^{t}\mathfrak{p}_{i})R_{P}\subseteq PR_{P}=0.$ Since $\min(-)$ is the minimal part of $\operatorname{Supp}(-)$, we proved that $\operatorname{Supp}(\bigcap_{i=1}^{t}\mathfrak{p}_{i})=\emptyset,$ and so $\bigcap_{i=1}^{t}\mathfrak{p}_{i}=0$. Therefore, $R$ is reduced. ∎
Notation 2.4.
By $(G_{i})$ (resp. $(R_{i})$) we mean $R_{\mathfrak{p}}$ is Gorenstein (resp. regular) for all $\mathfrak{p}\in\operatorname{Spec}(R)$ of height at most $i$.
Recall that a module $M$ satisfies $(S_{i})$ if $0pt(M_{\mathfrak{p}})\geq\min\{i,\dim(M_{\mathfrak{p}})\}$ for all $\mathfrak{p}\in\operatorname{Spec}(R)$.
Proposition 2.5.
Suppose that $\operatorname{id}_{R}(\mathfrak{p})<\infty$ for some ${\mathfrak{p}}\in\operatorname{Assh}(R)$. Then $R$ is quasi-reduced.
Proof.
Following Bass’ conjecture, $R$ is Cohen-Macaulay. In particular, $R$ satisfies Serre’s condition $(S_{1})$. Let ${\mathfrak{q}}\in\operatorname{Ass}(R)=\min(R)$. We have two possibilities i) $\mathfrak{q}=\mathfrak{p}$ or ii) $\mathfrak{q}\neq\mathfrak{p}$.
  1. i)
    In the first case, $R_{\mathfrak{q}}$ is a field. To see this, recall that $\operatorname{id}_{R_{\mathfrak{q}}}(\mathfrak{q}R_{\mathfrak{q}})=\operatorname{id}_{R_{\mathfrak{p}}}(\mathfrak{p}R_{\mathfrak{p}})$ is finite, because $\operatorname{id}_{R}(\mathfrak{p})<\infty$. This in turn implies that $R_{\mathfrak{q}}$ is zero-dimensional local and regular. From this, $R_{\mathfrak{q}}$ is a field.
  2. ii)
    In the second case, $R_{\mathfrak{q}}$ is Gorenstein. Indeed, we have $R_{\mathfrak{q}}=\mathfrak{p}R_{\mathfrak{q}}$. Consequently, $\operatorname{id}_{R_{\mathfrak{q}}}(R_{\mathfrak{q}})=\operatorname{id}_{R_{\mathfrak{q}}}(\mathfrak{p}R_{\mathfrak{q}})$, and recall that the later is finite, because $\mathfrak{p}$ is of finite injective dimension. By this, $R_{\mathfrak{q}}$ is Gorenstein.
In both cases, $R_{\mathfrak{q}}$ is Gorenstein. It remains to note that $(S_{1})$ and $(G_{0})$ imply the quasi-reduced condition. ∎
Proposition 2.6.
Suppose that $\operatorname{id}_{R}(\mathfrak{p})<\infty$ for all ${\mathfrak{p}}\in\operatorname{Spec}(R)$ of height at most one. Then $R$ is normal.
Proof.
Recall from the assumptions that $R$ satisfies Serre’s condition $(S_{2})$. Let ${\mathfrak{p}}\in\operatorname{Spec}(R)$ be of height at most one. Since $\operatorname{id}_{R}(\mathfrak{p})<\infty$, $\operatorname{id}_{R_{\mathfrak{p}}}(\mathfrak{p}R_{\mathfrak{p}})<\infty$. According to Auslander-Buchsbaum-Serre, $R_{\mathfrak{p}}$ is regular. So, $R$ satisfies $(R_{1})$. It remains to note that if a ring satisfies Serre’s condition $(S_{2})$ and $(R_{1})$ is normal. ∎
We cite [35] for basics of quasi-normal rings.
Proposition 2.7.
Suppose that $\operatorname{id}_{R}(\mathfrak{p})<\infty$ for some ${\mathfrak{p}}\in\operatorname{Spec}(R)$ of positive height. Then $R$ is quasi-normal.
Proof.
As Proposition 2.6, the ring $R$ satisfies $(S_{2})$. Let ${\mathfrak{q}}\in\operatorname{Spec}(R)$ be of height at most one. We may assume that $\mathfrak{p}$ in not maximal, otherwise the ring becomes regular. We have two possibilities i) $\mathfrak{q}\subseteq\mathfrak{p}$ or ii) $\mathfrak{q}\nsubseteq\mathfrak{p}$.
  1. i)
    We localize the ring at $\mathfrak{p}$ and by using the assumption $\operatorname{id}_{R}(\mathfrak{p})<\infty$, we know that $R_{\mathfrak{p}}$ is regular. Since $\mathfrak{q}\subseteq\mathfrak{p}$, we have $R_{\mathfrak{q}}\cong(R_{\mathfrak{p}})_{\mathfrak{q}R_{\mathfrak{p}}}$. Since localization of a regular ring is regular we deduce that $R_{\mathfrak{q}}$ is regular. So, $R$ satisfies $(R_{1})$.
  2. ii)
    Suppose $\mathfrak{q}\nsubseteq\mathfrak{p}$. We claim that $\mathfrak{p}\nsubseteq\mathfrak{q}$. Indeed, suppose not. Then $\mathfrak{p}\subseteq\mathfrak{q}$. Due to the assumption $\operatorname{ht}(\mathfrak{p})\geq 1$, and the fact that $\operatorname{ht}(\mathfrak{q})\leq 1$, we deduce that $\mathfrak{p}=\mathfrak{q}$. This contradicts the assumption from ii). So, $\mathfrak{p}\nsubseteq\mathfrak{q}$. This gives us the natural isomorphism $\mathfrak{p}R_{\mathfrak{q}}\cong R_{\mathfrak{q}}$. Since $\operatorname{id}_{R}(\mathfrak{p})<\infty$, its localizations are as well. In particular, $\operatorname{id}_{R_{\mathfrak{q}}}(R_{\mathfrak{q}})=\operatorname{id}_{R_{\mathfrak{q}}}(\mathfrak{p}R_{\mathfrak{q}})<\infty$. By definition $R_{\mathfrak{q}}$} is Gorenstein. So, $R$ satisfies $(G_{1})$.
In both cases $R$ is $(G_{1})$. It remains to note that if a ring satisfies $(S_{2})$ and $(G_{1})$, then it is quasi-normal. ∎
Corollary 2.8.
Suppose that $\operatorname{id}_{R}(\mathfrak{p})<\infty$ for some ${\mathfrak{p}}\in\operatorname{Spec}(R)$. Then $R$ is quasi-reduced.
Proof.
If $\mathfrak{p}$ is in $\operatorname{Assh}(R)$, then the desired claim is subject of Proposition 2.5. So, without loss of generality, we may assume that $\dim(R)>0$ and that $\mathfrak{p}\in\operatorname{Spec}(R)\setminus\operatorname{Assh}(R)$. By Bass’ conjecture, $R$ is Cohen-Macaulay. Since the ring is equi-dimensional we have $\operatorname{ht}(\mathfrak{p})>0$. It remains to apply Proposition 2.7. ∎
Proposition 2.9.
Let $(R,\mathfrak{m},k)$ be excellent $\operatorname{UFD}$ containing $k$ and of dimension >3>3. Suppose that $\operatorname{id}_{R}(\mathfrak{p})<\infty$ for all ${\mathfrak{p}}\in\operatorname{Spec}(R)$ of height two. Then $R$ is regular.
Proof.
This is straightforward to see $R$ satisfies Serre’s condition $(S_{3})$ and $(R_{2})$. Let $M:=\operatorname{Syz}_{2}(k)$. It is reflexive and locally free over the punctured spectrum. According to the Miller’s paper [29], there is a free module $F$ and a prime ideal $\mathfrak{p}$ of height at most two that fits in the following Bourbaki sequence
$0\longrightarrow F\longrightarrow M\longrightarrow\mathfrak{p}\longrightarrow 0\quad(\ast)$
Recall that $R$ is Cohen-Macaulay. Thanks to a result of Murthy [11, 3.3.19] $R$ is Gorenstein. This shows that $\operatorname{id}_{R}(F)<\infty$. We apply this in $(\ast)$ to see $\operatorname{id}_{R}(M)$ is finite. Thus, $\operatorname{id}_{R}(k)<\infty$. Thanks to Auslander-Buchsbaum-Serre, $R$ is regular. ∎
What about rings of mixed characteristic? More formally, we ask:
Problem 2.10.
How can find the mixed characteristic analogue of Bourbaki sequence and (related) prime ideals?***We are interested on it, even if the ring is regular.
If we allow more ideals are of finite injective dimension, then we present the following 3 positive answers:
Proposition 2.11.
Let $(R,\mathfrak{m})$ be local. Suppose that $\operatorname{id}_{R}(\mathfrak{a})<\infty$ for all ideals of big-height at most two. Then $R$ is regular.
Proof.
By Proposition 2.6 $R$ is normal. We assume that $\dim R>0$. Thanks to Bass, $R$ is Cohen-Macaulay. For any nonzero $x\in\mathfrak{m}$, $(xR)$ is height-unmixed and of height one. By assumption $\operatorname{id}_{R}(xR)$ is finite. Since $R$ is domain, $xR\cong R$. Thus, $\operatorname{id}_{R}(R)<\infty$. In other words, $R$ is Gorenstein. This allows us to assume that $\operatorname{pd}_{R}(\mathfrak{a})<\infty$ for all ideal $\mathfrak{a}$ of big-height at most two. This plus the normality and Cohen-Macaulay enable us to apply a result of Auslander [6, Corollary 5], and deduce that $R$ is regular. ∎
Proposition 2.12.
Let $(R,\mathfrak{m},k)$ be a complete $\operatorname{UFD}$. Suppose that $\operatorname{id}_{R}(\mathfrak{a})<\infty$ for all unmixed ideal of height two. Then $R$ is regular.
Proof.
We may assume that $\dim(R)>1$. Let $x,y$ be a parameter sequence. It is height unmixed. By assumption, $\operatorname{id}_{R}(x,y)<\infty$. Again, Bass’ conjecture says that the ring is Cohen-Macaulay. By the mentioned result of Murthy, $R$ is Gorenstein. So, $\operatorname{pd}_{R}(\mathfrak{a})<\infty$ for all unmixed ideal of height two. It remains to apply the discussion below [6, Corollary 5]. ∎
Here, is another variation of Proposition 2.11:
Observation 2.13.
Suppose any two generated ideal of $R$ is of finite injective dimension. Then $R$ is a Gorenstein $\operatorname{UFD}$.
Proof.
The ring $R:=(1)$ is of finite injective dimension. So, $R$ is Gorenstein. ***Alternatively, use [32, II.5.5]. Over Gorenstein rings, finite injective dimension implies finite projective dimension (see [11, 3.1.25]). Now, result is clear by [12, 5.3]. ∎
3. Modules with few generators
We start by the following easy exercise:
Fact 3.1.
Let $R$ be a local domain. Suppose there is a cyclically presented module $M$ of finite injective dimension. Then $R$ is Gorenstein.
Proof.
Let $M=R/xR$ for some $x$. Set $i:=\operatorname{id}_{R}(M)<\infty$ and look at $0\to R\stackrel{{\scriptstyle x}}{{\longrightarrow}}R\to M\to 0$. This gives $\operatorname{Ext}^{i+1}_{R}(k,R)\stackrel{{\scriptstyle x}}{{\longrightarrow}}\operatorname{Ext}^{i+1}_{R}(k,R)\to\operatorname{Ext}^{i+1}_{R}(k,M)=0$. By Nakayama, $\operatorname{Ext}^{i+1}_{R}(k,R)=0$, and so $\operatorname{id}_{R}(R)<\infty$. By definition, $R$ is Gorenstein. ∎
The following is nontrivial:
Fact 3.2.
(Peskine-Szpiro) Suppose there is a cyclic module $M$ of finite injective dimension. Then $R$ is Gorenstein.
Corollary 3.3.
Suppose $R$ is Gorenstein. If for some $\mathfrak{p}\in\operatorname{Spec}(R)$ one has $\operatorname{id}_{R}(\mathfrak{p})<\infty$, then $R$ is an integral domain.
Proof.
We have $\operatorname{pd}_{R}(\mathfrak{p})<\infty$. Then $R$ is an integral domain. ∎
Corollary 3.4.
Suppose for some principal ideal $\mathfrak{p}\in\operatorname{Spec}(R)$ one has $\operatorname{id}_{R}(\mathfrak{p})<\infty$. Then $R$ is an integral domain.
Proof.
Thanks to Fact 3.2, $R$ is Gorenstein. By Corollary 3.3, $R$ is an integral domain. ∎
Here, we present a new proof of [36, Theorem 5.2] and [31, Theorem 2.11(iv)].
Corollary 3.5.
Let $I\lhd R$ be nonzero. If $\operatorname{id}_{R}(R/I)<\infty$ then $I$ has an $R$-regular element.
Proof.
Thanks to Fact 3.2, $R$ is Gorenstein. By [11, 3.1.25], $\operatorname{pd}_{R}(R/I)<\infty$. Equivalently, II has a finite free resolution. By a famous result of Burch [11, Corollary 1.4.7] II contains a regular element. ∎
Notation 3.6.
By $\mu(-)$ we mean the minimal number of elements that needs to generates a finitely generated module ()(-).
In the final section, we extend the next result, by removing quasi-normality.
Proposition 3.7.
Let $(R,\mathfrak{m})$ be quasi-normal, and let $M$ be such that
  1. i)
    $\operatorname{id}_{R}(M)<\infty$,
  2. ii)
    $M$ is reflexive,
  3. iii)
    $\mu(M)\leq 2$.
Then $R$ is Gorenstein.
Proof.
By Fact 3.2 we may assume that $M$ is not cyclic. In the light of Bass’ conjecture, $R$ is Cohen-Macaulay. Let $d:=\dim(R)$, and $\underline{x}:=x_{1},\ldots,x_{d}$ be a system of parameter. In the light of [32, II.5.4] we have
$\operatorname{Ext}^{d}_{R}(\operatorname{Hom}_{R}(k,R/\underline{x}R),M)\cong k\otimes_{R}\frac{M}{\underline{x}M}\cong\frac{M}{\mathfrak{m}M}.$
The notation $r(R)$ stands for type of $R$. We deduce by this that $\oplus_{r(R)}\operatorname{Ext}^{d}_{R}(k,M)\cong k\oplus k.$ Let us denote $\mu_{i}(M)$ for the $i^{th}$ Bass number of $M$, i.e., the number of copies of $E_{R}(k)$ in the minimal injective resolution of $M$. Following this terminology, we deduce $\mu_{d}(M)r(R)=2$. In the case $r(R)=1$, there is nothing to proof, as the rings of type one are Gorenstein. So, without loss of generality, $r(R)=2$. This implies that $\mu_{d}(M)=1$. Since $M$ is torsion-free, it is of maximal dimension. Then, in view of [16, 3.4.c] we observe that $M$ is maximal Cohen-Macaulay. Without loss of generality we may and do assume that the ring is complete. Over Cohen-Macaulay and complete ring the canonical module exists, as it is unique, we denoted it by $\omega_{R}$. Since $M$ is maximal Cohen-Macaulay and of finite injective dimension, we use a result of Sharp [31] to present the identification $M=\oplus\omega_{R}$ (also, see [11, 3.3.28]). Since $\mu(M)=2$ and $\mu(\omega_{R})=r(R)=2$, we deduce that $M=\omega_{R}$. By $(-)^{\ast}$ we mean $\operatorname{Hom}_{R}(-,R)$. Thanks to [21, Appendix], there is an exact sequence
$0\longrightarrow\omega^{\ast}\longrightarrow R^{2}\longrightarrow\omega_{R}\longrightarrow 0\quad(\ast)$
Recall that $\operatorname{Ext}^{1}_{R}(R^{2},R)=0$, and apply $\operatorname{Hom}_{R}(-,R)$ to the displayed short exact sequence $(\ast)$, gives us the following:
[…]
From this we conclude that $\operatorname{Ext}^{1}_{R}(\omega_{R},R)=0$. Now, we are going to apply [2, Section 5.3] and to deduce that $R$ is Gorenstein. ∎
The above proof can be deduce from the following nice observation:
Observation 3.8.
Suppose $N$ is finitely generated and of finite injective dimension. Then $\mu(N)\geq\it{type}(R)$.
Proof.
The ring $R$ is Cohen-Macaulay, and we may assume it is complete. Now, we look at the maximal Cohen-Macaulay module $M:=R$. By [15, Corollary 3.15] there is surjection $M^{\oplus\mu(N)}\to\omega_{R}\to 0$. It remains to note from [11, 3.3.11(c)] that $\mu(\omega_{R})=\it{type}(R)$. ∎
Proposition 3.9.
Let $(R,\mathfrak{m})$ be quasi-normal of type at most two, and let $M$ be an indecomposable module such that
  1. i)
    $\operatorname{id}_{R}(M)<\infty$,
  2. ii)
    $M$ is reflexive,
  3. iii)
    $\mu(M)\leq 4$.
Then $R$ is Gorenstein.
Proof.
We may assume that $\mu(M)>3$. Recall from $i)$ that $R$ is Cohen-Macaulay. Suppose on the way of contradiction that $R$ is not Gorenstein. Then, we may assume that $r(R)=2$. By the above argument, $2\mu_{d}(M)=\mu_{d}(M)r(R)=\mu(M)$. From this, $\mu(M)$ is even. In fact $\mu(M)=4$ and so $\mu_{d}(M)=2$. Recall that $M$ is $(S_{2})$. According to [1, Theorem 3] we know $M$ is Cohen-Macaulay, and so maximal Cohen-Macaulay. Since it is of finite injective dimension, $M=\oplus\omega_{R}$. Since $M$ is indecomposable, $M=\omega_{R}$. But, $4=\mu(M)=\mu(\omega_{R})=r(R)=2.$ This contradiction completes the proof. ∎
Corollary 3.10.
Let $(R,\mathfrak{m})$ be quasi-normal of type two, and let $M$ be a module such that $\operatorname{id}_{R}(M)<\infty$. Then $\mu(M)$ is even.
Remark 3.11.
In the previous 3 results we may replace the quasi-normal assumption with the integral dimain assumption, see [2, Corollary 5.11].
The following is dual to [34, 6.2.4], and is clear from Bass’ conjecture:
Observation 3.12.
Suppose there is a finite length module $A$ of finite injective dimension. Then $R$ is Cohen-Macaulay.
Proof.
We may assume $R$ is complete. By Matlis, $M:=\operatorname{Hom}_{R}(A,E_{R}(k))$ is of finite length and of finite flat dimension. By [19], $\operatorname{pd}_{R}(M)<\infty$. Since $\ell(M\otimes R)<\infty$, and by intersection theorem, $\dim R\leq\operatorname{pd}_{R}(M)=0pt(R)-0pt(M)=0pt(R)$, as claimed. ∎
By the same reasoning, the following slightly extends [34, 6.2.4]:
Observation 3.13.
Suppose there is an artinian module $A$ of finite projective dimension. Then $R$ is Cohen-Macaulay.
By applying more advance results, we show:
Observation 3.14.
Suppose there is an $\mathfrak{m}$-torsion module $A$ of finite projective dimension. Then $R$ is Cohen-Macaulay.
Proof.
Since $A$ is $\mathfrak{m}$-torsion, $A^{\vee}$ is complete (see [30, 4.2]). In particular, there is an acyclic complex of injective modules resolves the complete module $A^{\vee}$. Recall by the work of Andre that $R$ is equipped with a big Cohen-Macaulay module. By [30, 7.6] $R$ is Cohen-Macaulay. ∎
The following drops the Cohen-Macaulay assumption from [36, Proposition 2.40]:
Corollary 3.15.
Let $(R,\mathfrak{m})$ be a 2-dimensional ring, $I$ an ideal of grade two and finite projective dimension. Then $r(R/I)=r(R)(\mu(I)-1)$.
Proof.
Recall from $2=\dim R\geq\operatorname{ht}(I)\geq\operatorname{grade}(I,R)=2$ that $I$ is $\mathfrak{m}$-primary. From $0\to I\to R\to R/I\to 0$ we deduce that the artinian module $R/I$ is of finite projective dimension. By Observation 3.14, $R$ is Cohen-Macaulay. The exact sequence $\ldots\to H^{i}_{\mathfrak{m}}(I)\to H^{i}_{\mathfrak{m}}(R)\to H^{i}_{\mathfrak{m}}(R/I)\to\ldots$ gives us $0pt(I)=1$. By the Cohen-Macaulay property and Auslander-Buchsbaum formula, $\operatorname{pd}_{R}(I)=0pt(R)-0pt(I)=1$. These allow us to apply [36, Proposition 2.40] and conclude $r(R/I)=r(R)(\mu(I)-1)$. ∎
Peskine-Szpiro [32, Corollary II.5.7] study the case $\mu(I)=2$. Here, is the case $\mu(I)=3$:
Corollary 3.16.
Let $(R,\mathfrak{m})$ be a 2-dimensional ring, $I$ an ideal of grade two and finite injective dimension. If $\mu(I)=3$, then:
  1. i)
    $R$ is Gorenstein,
  2. ii)
    $R/I$ is Cohen-Macaulay and of type two.
Proof.
i) Recall that $R$ is Cohen-Macaulay. Let $\underline{x}=x_{1},x_{2}$ be a parameter sequence. Then $\operatorname{Ext}^{2}_{R}(\operatorname{Hom}_{R}(k,R/\underline{x}R),I)\cong\frac{I}{\mathfrak{m}I}.$ This gives $r(R)\mu_{2}(I)=3$. Recall that $0pt(I)=1<2=\dim(I)$. Thanks to [16, Proposition 3.8] we know that $\mu_{2}(I)\geq 2$. Plugging this in the previous equality yields that $r(R)=1$, i.e., $R$ is Gorenstein.
ii) Since $R/I$ is zero-dimensional, it is Cohen-Macaulay. Due to the first item we know $R$ is Gorenstein. This yields that $\operatorname{pd}_{R}(I)<\infty$. By Corollary 3.15 $R/I$ is of type two. ∎
4. Auslander’s zero-divisor conjecture
As a basic tool, we prepare a version of Auslander’s zero-divisor conjecture for modules of finite injective dimension. This has its own importance.
Proposition 4.1.
Let $M$ be finitely generated, and of finite injective dimension. Then any $M$-regular element is $R$-regular.
Proof.
By [13, Proposition 3.6] it is enough to show that $\operatorname{grade}(\mathfrak{p},M)\leq\operatorname{grade}(\mathfrak{p},R)$ for all $\mathfrak{p}\in\operatorname{Spec}(R).$ Following Bass’ conjecture, $R$ is Cohen-Macaulay. Then
$\begin{array}[]{ll}\operatorname{grade}(\mathfrak{p},M)&\stackrel{{\scriptstyle(1)}}{{\leq}}0pt(M_{\mathfrak{p}})\\ &\stackrel{{\scriptstyle(2)}}{{\leq}}\dim(M_{\mathfrak{p}})\\ &\stackrel{{\scriptstyle(3)}}{{\leq}}\operatorname{id}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})\\ &\stackrel{{\scriptstyle(4)}}{{=}}0pt(R_{\mathfrak{p}})\\ &\stackrel{{\scriptstyle(5)}}{{=}}\operatorname{grade}(\mathfrak{p},R),\end{array}$
where 1) is in [26, Exercise 16.5]; 2) is in [11, Proposition 1.2.12]; 3) and 4) are in [11, Theorem 3.1.17]. For 5), see [11, Theorem 2.1.3(b)] and recall that $R$ is Cohen-Macaulay. ∎
Let us reconstruct Proposition 4.1 by a new argument:
Proposition 4.2.
Let $M$ be finitely generated, and of finite injective dimension. Then any $M$-regular element is $R$-regular.
Proof.
[…]
where
  1. 1):
    see [11, 1.2.27];
  2. 2):
    see [11, Theorem 3.3.5(b)].
Also, 3) is trivial. Since
[…]
Let us present a modern proof of Proposition 2.2:
Corollary 4.3.
Suppose $\min(R)$ is singleton. If for some $\mathfrak{p}\in\operatorname{Spec}(R)$ one has $\operatorname{id}_{R}(\mathfrak{p})<\infty$, then $R$ is an integral domain.
Proof.
We know $R$ is Cohen-Macaulay, and in particular, $\min(R)=\operatorname{Ass}(R)$. Let $\min(R)=\{P\}$. Let $\pi:R\to R_{\mathfrak{p}}$ be the natural localization map, sending $r$ to $r/1$. Let $r\in\ker(\pi)$. By definition, there is $x\in R\setminus\mathfrak{p}$ such that $rx=0$. We claim that $x$ is $\mathfrak{p}$-regular. Indeed, if not, then
$x\in\operatorname{Zd}(\mathfrak{p})=\cup_{\mathfrak{q}\in\operatorname{Ass}(\mathfrak{p})}\mathfrak{q}\subseteq\cup_{\mathfrak{q}\in\operatorname{Ass}(R)}\mathfrak{q}=P.$
Since $\min(R)=\{P\}$, we see $P\subseteq\mathfrak{p}$. So, $x\in\mathfrak{p}$. This contradiction shows that $x$ is $\mathfrak{p}$-sequence. Now, we apply the assumption along with Proposition 4.2 to observe that $x$ is $R$-sequence. From this $r=0$, i.e., $\pi$ is injective. Now recall from $\operatorname{id}_{R}(\mathfrak{p})<\infty$ that $R_{\mathfrak{p}}$ is regular and local. Regular local rings are domain (see [11, Proposition 2.2.3]). From this, $R_{\mathfrak{p}}$ is an integral domain. Since $R\subseteq R_{\mathfrak{p}}$ we get to the desired claim. ∎
Let us reprove Auslander’s zero-divisor conjecture, over Cohen-Macaulay rings:
Observation 4.4.
Let $(R,\mathfrak{m})$ be Cohen-Macaulay, $M$ be finitely generated with $\operatorname{pd}_{R}(M)<\infty$. If $x$ is $M$-regular, then $x$ is $R$-regular.
Proof.
We may and do assume that $R$ is complete. In particular, $\omega_{R}$ exists. The assumption $\operatorname{pd}_{R}(M)<\infty$ implies that $\operatorname{id}_{R}(M\otimes_{R}\omega_{R})<\infty$. We proved in the previous proposition that $\operatorname{Ass}(M\otimes_{R}\omega_{R})=\operatorname{Ass}(\operatorname{Hom}_{R}(\omega_{R},M\otimes_{R}\omega_{R}))$. We are going to use [31, Theorem 2.9] to deduce that $\operatorname{Hom}_{R}(\omega_{R},M\otimes_{R}\omega_{R})\cong M$. Combining this along with the previous observation, we know that $\operatorname{Ass}(M\otimes_{R}\omega_{R})=\operatorname{Ass}(M)$. So, $x$ is $(M\otimes_{R}\omega_{R})$-regular. Thanks to Proposition 4.1, $x$ is $R$-regular, as claimed. ∎
Remark 4.5.
i) Recall that [4] asks the computation of associated prime ideals of tensor product, as a sample, we proved in the setting of Observation 4.4 that $\operatorname{Ass}(M\otimes_{R}\omega_{R})=\operatorname{Ass}(M).$ In particular, we reproved [4, Corollary 4.6] by a different argument.
ii) By symmetry, and in the proof of Observation 4.4, one may use [4, Corollary 4.6] instead of [31, Theorem 2.9].
Recall from [8] module $T$ is said to be tor-rigid if there is a non-negative integer $n$ such that for every finitely generated $R$-module $M$, vanishing of $\operatorname{Tor}_{n}^{R}(T,M)$ implies $\operatorname{Tor}_{n+i}^{R}(T,M)$ vanishes for all $i\geq 0$. By Auslander [8, 4.3], this implies the zero-divisor property.
Remark 4.6.
Despite to above useful duality between modules of finite projective dimension and modules of finite injective dimension, let us present a different situation. It is easy to see modules of projective dimension one are tor-rigid. But modules of injective dimension one are not necessarily tor-rigid. This may appear even over 1-dimensional integral domains of type two. For instance, let $R:=k[[x^{3},x^{4},x^{5}]]$. Then $(x^{3},x^{4})$ is of injective dimension one, but not tor-rigid.
For more connections to these topics, we cite [4].
5. Almost complete-intersection
In this section $(S,\mathfrak{n})$ is regular, $I\lhd S$ is an ideal, and $R:=S/I$. We study the following:
Question 5.1.
(See [20]) Suppose $\operatorname{id}_{R}(I/I^{2})$ is finite. Is $I$ generated by a regular sequence?
Here, we support it by the following two observations:
Observation 5.2.
If $R$ is zero-dimensional, then $I$ is generated by a regular sequence.
Proof.
The $R$-module $I/I^{2}$ is injective, as injective dimension is bounded by depth. By Matlis decomposition, $I/I^{2}=\oplus\omega_{R}$. Now, let $n\geq 0$ be such that $\mathfrak{m}^{n}\neq 0$ but $\mathfrak{m}^{n+1}=0$. Then
$\mathfrak{m}^{n}(I+I^{2})=(\mathfrak{m}^{n}I+I^{2})\subseteq\mathfrak{m}^{n+1}+I^{2}=0.$
From this, $\mathfrak{m}^{n}\omega_{R}=0.$ As $\omega_{R}$ is faithful, we conclude that $n=0$. In other words, $R$ is field. So, $I$ is generated by a regular sequence. ∎
Corollary 5.3.
Question 5.1 reduces to locally complete-intersection ideals over the punctured spectrum.
Proof.
We proceed by induction on $d:=\dim R$. The case $d=0$ is subject of Observation 5.2. So, we may assume $d>0$ and suppose the desired claim holds for rings of dimension <d<d. The assumptions of Question 5.1 behave well with respect to localization for all $\mathfrak{q}\in\operatorname{Var}(I)\setminus\{\mathfrak{n}\}$, and note that $\dim(R_{\mathfrak{q}})<d$. These allow us to apply the inductive step and assume in addition that $I$ is locally complete-intersection over the punctured spectrum. This completes the proof. ∎
Observation 5.4.
Suppose $I$ is prime, $\mu(I)\leq\operatorname{ht}(I)+1$ and $\operatorname{id}_{R}(I/I^{2})<\infty$. Then $I$ is generated by a regular sequence.
Proof.
By Bass’ conjecture, $R$ is Cohen-Macaulay. Then $K_{R}=\omega_{R}$ is of finite injective dimension. Suppose on the way of contradiction that $\mu(I)=\operatorname{ht}(I)+1$. In the light of [28, Proposition 1] there is an exact sequence
$0\longrightarrow K_{R}\longrightarrow R^{n}\longrightarrow I/I^{2}\longrightarrow 0,$
which show that $\operatorname{id}_{R}(R)<\infty$. Over Gorenstein rings, finite injective dimension implies finite projective dimension. So, $\operatorname{pd}_{R}(I/I^{2})$ is finite. Follows from [37] that $I$ is generated by a regular sequence, as claimed. ∎
Let us present a new proof of [25]:
Theorem 5.5.
(Kunz) Almost complete intersection domains are not Gorenstein rings.
Proof.
We adopt the previous notation. Suppose on the way of contradiction that $R$ is Gorenstein. Then from $0\to K_{R}\to R^{n}\to I/I^{2}\to 0,$ and finiteness of $\operatorname{id}_{R}(R)$, we conclude that $\operatorname{id}_{R}(I/I^{2})<\infty$. By the previous result, $\operatorname{ht}(I)=\mu(I)$, a contradiction. ∎
Corollary 5.6.
Suppose $I$ is prime and generated by $3$ elements. If $\operatorname{id}_{R}(I/I^{2})<\infty$, then $I$ is generated by a regular sequence.
Proof.
We have $\operatorname{ht}(I)\leq 3$. In the case $\operatorname{ht}(I)=1$ there is nothing to prove, because it becomes principal as $S$ is $\operatorname{UFD}$. In the case $\operatorname{ht}(I)=3$ the claim is in [11, 1.2.21]. Then we may assume that $\operatorname{ht}(I)=2$. The desired claim follows by Observation 5.4. ∎
Observation 5.7.
If $R$ is Gorenstein and $\operatorname{id}_{R}(I/I^{2})<\infty$, then $I$ is generated by a regular sequence.
Proof.
Since $R$ is Gorenstein, $\operatorname{pd}_{R}(I/I^{2})<\infty$. In view of [9, main result] $I$ generated by a regular sequence. ∎
For more connections to [20], we cite [3].
6. Matlis and canonical module
A ring $R$ of dimension $d:=\dim(R)$ is called quasi-Gorenstein, if $H^{d}_{\mathfrak{m}}(R)\cong E_{R}(k)$. When is $H^{d}_{\mathfrak{m}}(R)\twoheadrightarrow E_{R}(k)$ surjective? The following may regard as a higher version of [27, 15.17]:
Proposition 6.1.
Suppose $R$ is complete, generically Gorenstein and Cohen-Macaulay. Then $H^{\dim R}_{\mathfrak{m}}(R)\twoheadrightarrow E_{R}(k)$ is surjective.
Proof.
Let $d:=\dim(R)$. Since the ring is complete and Cohen-Macaulay, we know $\omega_{R}$ exists. Since $R$ is generically Gorenstein, $\omega_{R}\lhd R$. Suppose first that $d:=\dim R=0$. Then $R$ is Gorenstein, and so $\omega_{R}=E_{R}(k)=R$. To see $H^{d}_{\mathfrak{m}}(R)\stackrel{{\scriptstyle}}{{\twoheadrightarrow}}E_{R}(k)$, it remains to note that $H^{d}_{\mathfrak{m}}(R)=R$. So, we may assume that $d>0$. Either $\omega_{R}=R$, or $\omega_{R}\lhd R$ is of height one. Without loss of the generality we may assume that $\omega_{R}\lhd R$ is of height one. The Cohen-Macaulay assumption says $1=\operatorname{ht}(\omega_{R})=\operatorname{grade}(\omega_{R},R)$. Let $x\in\omega_{R}$ be a regular element. This gives an embedding $R\hookrightarrow Rx\subseteq\omega_{R}$. Apply $\operatorname{Hom}_{R}(-,E_{R}(k))$ to it, and use local duality, we observe that
[…]
In other words, the induced map $H^{d}_{\mathfrak{m}}(R)\twoheadrightarrow E_{R}(k)$ is surjective, as claimed. ∎
Proposition 6.2.
$d:=\dim(R)$. Suppose $H^{d}_{\mathfrak{m}}(R)\twoheadrightarrow E_{R}(k)$. If $d=0$, then $R$ is Gorenstein.
Proof.
Recall that
[…]
Let $I:=\ker(f)$. So, $R/I=E_{R}(k)$. Recall that $R\stackrel{{\scriptstyle\subseteq}}{{\longrightarrow}}E(R)=\oplus E_{R}(k)=\oplus R/I.$ From this, $I=0$. Thus, $R=E_{R}(k)$. Consequently, $R$ is Gorenstein. ∎
Let $R$ be a 1-dimensional integral domain and let $Q$ be the fraction field. Matlis proved that $Q/\omega_{R}=E_{R}(k)$. Here, is the higher version (recall that $Q/\omega_{R}=H^{1}_{\mathfrak{m}}(\omega_{R})$):
Proposition 6.3.
Suppose $R$ is complete, integral domain and Cohen-Macaulay. Let $I$ be a nonzero Cohen-Macaulay ideal. Let $d:=\dim(R)$. The following are equivalent:
  1. i)
    $\operatorname{id}_{R}(I)<\infty$,
  2. ii)
    $I=\omega_{R}$,
  3. iii)
    $H^{d}_{\mathfrak{m}}(I)=E_{R}(k)$.
Proof.
$i)\Rightarrow ii)$: Since $\dim I=d$, the ideal $I$ becomes maximal Cohen-Macaulay. The desired claim is in [11, 3.3.28].
$ii)\Rightarrow iii)$: Due to the right exactness, $H^{d}_{\mathfrak{m}}(\omega_{R})=H^{d}_{\mathfrak{m}}(R)\otimes_{R}\omega_{R}$. Recall that $H^{d}_{\mathfrak{m}}(\omega_{R})$ is artinian, and apply Matlis duality to it,
$\begin{array}[]{ll}H^{d}_{\mathfrak{m}}(\omega_{R})\stackrel{{\scriptstyle}}{{=}}H^{d}_{\mathfrak{m}}(\omega_{R})^{\vee\vee}&=\operatorname{Hom}_{R}\big{(}H^{d}_{\mathfrak{m}}(R)\otimes_{R}\omega_{R},E_{R}(k)\big{)}^{\vee}\\ &=\operatorname{Hom}_{R}\big{(}\omega_{R},\operatorname{Hom}_{R}(H^{d}_{\mathfrak{m}}(R),E_{R}(k)\big{)}^{\vee}\\ &=\operatorname{Hom}_{R}(\omega_{R},\omega_{R})^{\vee}=R^{\vee}=E_{R}(k).\end{array}$
$iii)\Rightarrow i)$: Let $D(-):=\operatorname{Hom}_{R}(-,\omega_{R})$. We apply local duality along with the assumption to deduce that $D(I)=H^{d}_{\mathfrak{m}}(I)^{\vee}=E_{R}(k)^{\vee}=\widehat{R}=R.$ Over maximal Cohen-Macaulay, $D^{2}(-)=\emph{Id}(-)$. Then $I\cong D^{2}(I)\cong D(R)\cong\omega_{R}.$ Since $\omega_{R}$ is of finite injective dimension, $I$ is as well. ∎
Corollary 6.4.
Let $R$ be a 1-dimensional integral domain and let $Q$ be the fraction field. Let $I\lhd R$ be nonzero. Then $Q/I=E_{R}(k)$ iff $I=\omega_{R}$.
The following was proved by Matlis in 1-dimensional complete Gorenstein domain case, and by Auslander [5] in the $\mathfrak{m}$-adic complete case. Since being complete in $\mathfrak{m}$-adic topology implies completeness in $R$-topology, so it extends both:
Observation 6.5.
Let $R$ be a quasi-local domain complete in $R$-topology. Then $\operatorname{Ext}^{1}_{R}(Q/R,R)\cong R$.
Proof.
Completeness in $R$-topology implies $\operatorname{Ext}^{1}_{R}(Q,R)=0$. Also, $\operatorname{Hom}_{R}(Q,R)=0$. The sequence $0\to R\to Q\to Q/R\to 0$ implies
$0=\operatorname{Hom}_{R}(Q,R)\to\operatorname{Hom}_{R}(R,R)\to\operatorname{Ext}^{1}_{R}(Q/R,R)\to\operatorname{Ext}^{1}_{R}(Q,R)=0,$
i.e., the claim follows.∎
Noetherian local rings are reduced***Sorry, we used “reduced” for different concepts. by Krull’s intersection theorem. So, the following is converse to Auslander [5]:
Observation 6.6.
Let $R$ be a quasi-local reduced domain such that $\operatorname{Ext}^{1}_{R}(Q/R,R)\cong R$. Then $R$ is complete in $R$-topology.
Proof.
The sequence $0\to R\to Q\to Q/R\to 0$ implies
$0=\operatorname{Hom}_{R}(Q,R)\to\operatorname{Hom}_{R}(R,R)\to\operatorname{Ext}^{1}_{R}(Q/R,R)\to\operatorname{Ext}^{1}_{R}(Q,R)\to\operatorname{Ext}^{1}_{R}(R,R)=0.$
[…]
7. Reflexivity and injective dimension
We start by recalling the following result:
Fact 7.1.
(Auslander, see [32, I. Proposition 4.14]). Suppose $M$ is of finite projective dimension and there is a $\mathfrak{p}\in\operatorname{Supp}(M)\cap\operatorname{Ass}(R)$. Then $\operatorname{Supp}(M)=\operatorname{Spec}(R)$.
Lemma 7.2.
Suppose $R$ is complete and there is a torsion-less module $M$ of finite injective dimension. Then $\operatorname{Supp}(M)=\operatorname{Spec}(R)$.
Proof.
Let $\tilde{M}:=\operatorname{Hom}_{R}(\omega_{R},M)$. Then
$\operatorname{Ass}_{R}(\tilde{M})=\operatorname{Supp}(\omega_{R})\cap\operatorname{Ass}(M)=\operatorname{Spec}(R)\cap\operatorname{Ass}(M)=\operatorname{Ass}(M).$
Since $M$ is torsion-less, then $M\subseteq R^{n}$ and so $\operatorname{Ass}_{R}(\tilde{M})=\operatorname{Ass}(M)\subseteq\operatorname{Ass}(R).$Recall that $\tilde{M}$ is of finite projective dimension. By using Fact 7.1 we know $\operatorname{Supp}(\tilde{M})=\operatorname{Spec}(R)$. Consequently, $\operatorname{Ass}(R)=\operatorname{Ass}_{R}(\tilde{M})=\operatorname{Ass}(M)\subseteq\operatorname{Ass}(R)$. The claim follows. ∎
The following extends Corollary 2.8.
Lemma 7.3.
Suppose there is a nonzero torsion-less module $M$ of finite injective dimension. Then $\widehat{R}$ is generically Gorenstein.
Proof.
We may assume that $R$ is complete. The ring $R$ is unmixed. Let $\mathfrak{p}\in\operatorname{Ass}(R)=\min{R}$. According to the previous lemma, we know $\mathfrak{p}\in\operatorname{Supp}(M)$. This shows that $M_{\mathfrak{p}}$ is nonzero (of finite) injective (dimension) over artinian ring $R_{\mathfrak{p}}$. As $\operatorname{id}(-)$ is bounded by depth, we deduce that $M_{\mathfrak{p}}=\oplus_{n>0}\omega_{R_{\mathfrak{p}}}$. Localization of torsionless is again torsionless. It follows that the canonical module of $R_{\mathfrak{p}}$ is torsion-less. This implies that $R_{\mathfrak{p}}$ is Gorenstein (see [3, Corollary 3.7]). ∎
The following is the main result of this section.
Proposition 7.4.
Suppose there is a nonzero reflexive module $L$ of finite injective dimension. Then $\widehat{R}$ is quasi-normal.
Proof.
We may assume that $R$ is complete. Let $\mathfrak{p}\in\operatorname{Spec}(R)$ be of height one. By previous lemma, $\mathfrak{p}\in\operatorname{Supp}(L)$. Set $A:=R_{\mathfrak{p}}$ and $M:=L_{\mathfrak{p}}$. Then $\operatorname{id}_{A}(M)=0pt_{A}(A)=1$. Consequently, $\operatorname{Ext}_{R}^{2}(\operatorname{Tr}(M^{\ast}),(M^{\ast})^{\ast})=0,$ where $\operatorname{Tr}(-)$ denotes the Auslander’s transpose (see [7] for its definition). Then the natural map
$f:((M^{\ast})^{\ast})\otimes_{R}(M^{\ast})\longrightarrow\operatorname{Hom}_{R}((M^{\ast})^{\ast},(M^{\ast})^{\ast})$
is surjective, as the cokernel of $f$ is $\operatorname{Ext}_{R}^{2}(\operatorname{Tr}(M^{\ast}),(M^{\ast})^{\ast})=0,$ see [7, Proposition (2.6)(a)]. It turns out that $M^{\ast}$ is free as an $A$-module. Since $M$ is reflexive, $M\cong M^{\ast\ast}$ is free as well. Thus, $\operatorname{id}(A)$ is finite. This means that $A=R_{\mathfrak{p}}$ is Gorenstein. So, $R$ satisfies $(G_{1})$. Following Bass’ conjecture, $R$ is Cohen-Macaulay. It remains to note that if a ring satisfies $(S_{2})$ and $(G_{1})$, then it is quasi-normal. ∎
Corollary 7.5.
Let $M$ be reflexive such that $\operatorname{id}_{R}(M)<\infty$, and $\mu(M)\leq 2$. Then $R$ is Gorenstein.
Proof.
Without loss of generality, we may assume $(R,\mathfrak{m})$ is complete in $\mathfrak{m}$-adic topology. In view of above proposition, the ring is quasi-normal. Now, apply Proposition 3.9. ∎
Corollary 7.6.
Let $(R,\mathfrak{m})$ be of type at most two, and let $M$ be an indecomposable reflexive module such that $\operatorname{id}_{R}(M)<\infty$, and $\mu(M)\leq 4$. Then $R$ is Gorenstein.
Corollary 7.7.
Let $(R,\mathfrak{m})$ be of type two, and let $M$ be a module such that $\operatorname{id}_{R}(M)<\infty$. Then $\mu(M)$ is even.
An $R$-module $M$ is called $k$-torsionless if $\operatorname{Ext}_{R}^{i}(\operatorname{Tr}M,R)=0$ for all $i\leq k$. Recall that torsionless is 1-torstionless and reflexive is 2-torstionless. To close the paper, we pose the following question and observation, inspired by the main result of this section, and present
Question 7.8.
Let $k\geq 0$. Suppose there is a nonzero $(k+1)$-torsionless module of finite injective dimension. When is $R$ satisfy the property $(G_{k})$?
Observation 7.9.
Let $A$ be a homomorphic image of Gorenstein ring. Suppose a nonzero module $N$ is $(k+2)$-torsionless and is of finite injective dimension. Then $A$ satisfies the property $(G_{k})$.
Proof.
Let $\mathfrak{p}\in\operatorname{Spec}(A)$ be of height at most $k$. Let $R:=A_{\mathfrak{p}}$ and $M:=N_{\mathfrak{p}}$. Then things reduced to show that $R$ is Gorenstein. The ring has canonical module. This helps to apply Lemma 7.2, and thus, $M$ is nonzero. Then, without loss of generality we can assume in addition that $k=d:=\dim R$, since $(\ell)$-torsionless behaves well with respect to localization. By definition, we know that $M$ is reflexive and $\operatorname{Ext}_{R}^{i}(M^{\ast},R)=0$ for all $i\leq d$, as $M^{\ast}$ is the second syzygy of $\operatorname{Tr}M$. Thanks to Bass, $d=\dim R=0pt(R)$. We proceed by induction on $d$ to deduce that $M$ is free. The initial case $d=0$ is clear from Lemma 7.3. By using induction hypothesis, we may assume in addition that $M$ is locally free over the punctured spectrum. Let $L:=M^{\ast}$. Then $L$ is locally free over the punctured spectrum and $\operatorname{id}_{R}(L^{\ast})\leq d$. So, $\operatorname{Ext}_{R}^{d+1}(\operatorname{Tr}L,L^{\ast})=0$ and that $\operatorname{Ext}_{R}^{i}(L,R)=0$ for all $i\leq d$. Under these assumptions, Kimura proved in [24, Theorem 2.4] that $L$ is free. Then $M\cong M^{\ast\ast}\cong L^{\ast}$ is free as well. Thus, $\operatorname{id}_{R}(R)\leq d$. Consequently, $R$ is Gorenstein. ∎
Corollary 7.10.
Suppose a nonzero module $M$ is $(d+2)$-torsionless and of finite injective dimension, where $d=\dim R$. Then $R$ is Gorenstein.
Proof.
We may assume the ring is complete. So, the result follows by Observation 7.9.∎
One can prove the case $d=0$, even without any use of Observation 7.9:
Fact 7.11.
Let $R$ artinian. Suppose a nonzero reflexive module $M$ of finite injective dimension. Then $R$ is Gorenstein.
Proof.
We dual $R^{n}\to M^{\ast}\to 0$. Then, it follows by definition that $M\cong M^{\ast\ast}\oplus X=R^{n}$ for some $n>0$ and $X$. Now, by Krull-Remak-Schmidt, $R$ is injective and so $R$ is Gorenstein.∎
8. On the injective dimension of Hom
This section is about:
Question 8.1.
(See [18, Question 2.18]). If there exists a nonzero $R$-module $M$ such that $\operatorname{Hom}_{R}(M,M)$ has finite injective dimension, then is $R$ Gorenstein?
Question 8.2.
(See [17, Question 4.9]). Let $M$ be a nonzero RR–module such that the injective dimensions of M and $\operatorname{Ext}^{i}_{R}(M,M)$ are finite for all $0\leq i\leq\dim(R)-0pt(M)$. Is $R$ Gorenstein?
8.1. Injective dimension of Hom
We start by the following easy exercise:
Fact 8.3.
Question 8.1 is true when $R$ is artinian.
Proof.
We may and do assume $M$ is indecomposable. Apply $\operatorname{Hom}_{R}(-,M)$ to $0\to\operatorname{Syz}(M)\to R^{n}\to M\to 0$ and obtain $0\to\operatorname{Hom}_{R}(M,M)\to\operatorname{Hom}_{R}(R^{n},M)\to C\to 0$, where $C$ is the cokernel. But, $\dim R=0$, and so $\operatorname{Hom}_{R}(M,M)$ is injective, e.g., $M^{n}=\operatorname{Hom}_{R}(M,M)\oplus C$. By Krull-Remek-Schmidit, $M^{m}=\operatorname{Hom}_{R}(M,M)$ for some $m\leq n$. Hence, $M=E_{R}(k)$, and consequentely, $R=\operatorname{Hom}_{R}(E_{R}(k),E_{R}(k))$ is injective. So, $R$ is Gorenstein. ∎
Here, we present a modern proof the above fact:
Proof.
One has $\operatorname{Ext}^{1}_{R}(M,\operatorname{Hom}(M,M))=0$, and so $M$ is free. This gives the injectivity of $R$. So, $R$ is Gorenstein. ∎
Proposition 8.4.
Let $M$ be a nonzero finite-length RR–module such that the injective dimension of $M$ and $\operatorname{Hom}_{R}(M,M)$ are finite. Then $R$ is Gorenstein.
Proof.
The ring is Cohen-Macaulay. Without loss of generality we may assume the ring is complete. Then we use Matlis duality. It turns out that $\operatorname{pd}(\operatorname{Hom}_{R}(M,M)^{\vee})<\infty$. But, $\operatorname{Hom}_{R}(M,M)^{\vee}=M\otimes M^{\vee}$. In view of [15, A.9(ii)], $R$ is Gorenstein. ∎
Proposition 8.5.
Let $M\subseteq E_{R}(k)$ be a nonzero finite-length RR–module such that the injective dimension of $M$ is finite. Then $R$ is Gorenstein.
Proof.
The ring is Cohen-Macaulay. Without loss of generality we may assume the ring is complete. Dualizing the embedding $0\to M\stackrel{{\scriptstyle\rho}}{{\longrightarrow}}E_{R}(k)$ yields the surjection $R\cong\widehat{R}\stackrel{{\scriptstyle\rho^{\vee}}}{{\longrightarrow}}M^{\vee}\to 0$. Let $I:=\ker(\rho^{\vee})$. Since $M^{\vee}=R/I$ we deduce that $I=\operatorname{Ann}(M^{\vee})=\operatorname{Ann}(M)$. Recall from $\operatorname{Hom}_{R}(M,M)^{\vee}=M\otimes M^{\vee}$ that
$M=\frac{M}{IM}=M\otimes M^{\vee}=\operatorname{Hom}_{R}(M,M)^{\vee}$
has finite projective dimension. Thus $M^{\vee}=R/I$ has finite injective dimension. Since it is cyclic, we apply the mentioned result of Peskine-Szpiro to deduce that $R$ is Gorenstein. ∎
The following was proved in [17] by spectral sequence argument:
Proposition 8.6.
Let $N$ be torsion-free with $0pt(N)\geq\dim R-1$ such that the injective dimension of both $N$ and $\operatorname{Ext}^{\ast}_{R}(M,N)$ are finite. Then $\operatorname{pd}(M)<\infty$.
Proof.
Without loss of generality we may $0pt(N)=\dim R-1$. By going to syzygies, we may assume that $M$ is maximal Cohen-Macaulay. Also, we assume in addition that $R$ is generically Gorenstein. We look at
$0\longrightarrow\omega_{R}^{n}\longrightarrow\omega_{R}^{m}\longrightarrow N\longrightarrow 0\quad(\ast)$
and recall that $\operatorname{Ext}^{1}_{R}(M,\omega_{R})=0$. This gives us,
$0\to\operatorname{Hom}_{R}(M,\omega_{R}^{n})\to\operatorname{Hom}_{R}(M,\omega_{R}^{m})\to\operatorname{Hom}_{R}(M,N)\to\operatorname{Ext}^{1}_{R}(M,\omega_{R})=0$
Since $\operatorname{id}(\operatorname{Hom}_{R}(M,N))<\infty$ we deduce
$\operatorname{Ext}^{\gg 0}_{R}(-,\operatorname{Hom}_{R}(M,\omega_{R}))^{n}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\operatorname{Ext}^{\gg 0}_{R}(-,\operatorname{Hom}_{R}(M,\omega_{R}))^{m}$
Now, if $\operatorname{Ext}^{\gg 0}_{R}(-,\operatorname{Hom}_{R}(M,\omega_{R}))\neq 0$, it then turns out that $n=m$, but this contradicts the rank-consideration in $(\ast)$. Thus, $\operatorname{Ext}^{\gg 0}_{R}(-,\operatorname{Hom}_{R}(M,\omega_{R}))=0$. By definition, $\operatorname{id}(\operatorname{Hom}_{R}(M,\omega_{R}))<\infty$. Since $M$ is maximal Cohen-Macaulay, $\operatorname{Hom}_{R}(M,\omega_{R})$ is as well. So, $\operatorname{Hom}_{R}(M,\omega_{R})=\oplus\omega_{R}.$ Thus, $M=M^{\dagger\dagger}=(\oplus\omega_{R})^{\dagger}=\oplus R$, as claimed. ∎
The ideal case has an affirmative answer:
Proposition 8.7.
Let $I$ be a nonzero ideal of $R$ such that the injective dimension of $\operatorname{Hom}_{R}(I,I)$ is finite. Then $R$ is Gorenstein.
Proof.
Let $d=\dim R$. We proceed by induction on $d$. The case $d=0$ is trivial. Suppose $d=1$. In this case are able to apply [18, 2.3] to deduce that $I=\Gamma_{\mathfrak{m}}(I)\oplus R^{m}$ for some $m\geq 0$. The ring is Cohen-Macaulay and $0pt(R)=d>0$. Thus, $\Gamma_{\mathfrak{m}}(R)=0$. But, $\Gamma_{\mathfrak{m}}(I)\subseteq\Gamma_{\mathfrak{m}}(R)=0$. Then $I\cong R^{m}$, and by rank consideration, $I\cong R$. Consequently, the injective dimension $R\cong\operatorname{Hom}_{R}(I,I)$ is finite. So, $R$ is Gorenstein. Therefore, we may assume $d>1$, and apply the inductive hypothesis to deduce that the natural map $\pi:R\longrightarrow\operatorname{Hom}_{R}(I,I)$ is an isomorphism over the punctured spectrum. Recall that $\pi$ is 1-to-1, as $I$ is torsion-free. From
$\zeta:=0\longrightarrow R\longrightarrow\operatorname{Hom}_{R}(I,I)\longrightarrow C:=\operatorname{Coker}(\pi)\longrightarrow 0,$
we deduce that $\operatorname{Supp}(C)\subseteq\{\mathfrak{m}\}$. Suppose $C$ is nonzero. Then
$\inf\{i:\operatorname{Ext}^{i}_{R}(C,R)\neq 0\}=0pt(R)=d>1.$
This implies that $\zeta\in\operatorname{Ext}^{1}_{R}(C,R)=0$, and so the sequence splits. Then $\operatorname{Hom}_{R}(I,I)=R\oplus C$. Since $C\subseteq\Gamma_{\mathfrak{m}}(I)\subseteq\Gamma_{\mathfrak{m}}(R)=0$ we see $C=0$. Then $\operatorname{Hom}_{R}(I,I)\cong R$. Consequently, the injective dimension $R\cong\operatorname{Hom}_{R}(I,I)$ is finite. So, $R$ is Gorenstein. ∎
Proposition 8.8.
Let $\mathbb{Q}\subseteq R$ and $M$ be torsion-free such that the injective dimension of $\operatorname{Hom}_{R}(M,M)$ is finite. Then $R$ is Gorenstein.
Proof.
Recall that $\operatorname{Supp}(M)=\operatorname{Spec}(R)$ any by the previous proof $M$ is locally-free in codimenstion one. Recall That ring is Cohen-Macaulay, and so $(S_{2})$. By [23, 5.5] we know $M$ has constant rank, since $\mathbb{Q}\subseteq R$, the rank is unit in $R$. By [18, 3.1] $R$ is Gorenstein. ∎
What about of big Cohen-Macaulay modules of finite injective dimension?
8.2. Remarks on the prime characteristic case
Rings in this subsection are of prime characteristic $p$ where $\operatorname{F}:R\to R$ is the Frobenius map. Each iteration of $\operatorname{F}_{n}$ of $\operatorname{F}$ defines a new $R$-module structure on the set $R$, and this $R$-module is denoted by ${{}^{\operatorname{F}_{n}}\!}R$, where $a\cdot b=a^{p^{n}}b$ for $a,b\in R$. Recall that $R$ is said to be $\operatorname{F}$-finite if for some (or equivalently, all) $n$, the algebra ${{}^{\operatorname{F}_{n}}\!}R$ is finitely generated as an $R$-module.
Observation 8.9.
Let $R$ be $\operatorname{F}-\emph{finite}$ and $\operatorname{F}-\emph{pure}$. If $\operatorname{id}_{R}(\operatorname{Hom}_{R}({{}^{\operatorname{F}_{n}}\!}R,{{}^{\operatorname{F}_{n}}\!}R))<\infty$, then $R$ is regular.
Proof.
Since $R$ is $\operatorname{F}-\emph{pure}$, then $R$ is a directed summand of ${{}^{\operatorname{F}_{n}}\!}R$. This shows ${{}^{\operatorname{F}_{n}}\!}R$ is a direct summand of $\operatorname{Hom}_{R}({{}^{\operatorname{F}_{n}}\!}R,{{}^{\operatorname{F}_{n}}\!}R)$. Then $\operatorname{id}({{}^{\operatorname{F}_{n}}\!}R)<\infty$. Since the ring is Gorenstein by [18, Prop 3.2], $\operatorname{pd}({{}^{\operatorname{F}_{n}}\!}R)<\infty$. By a result of Kunz, $R$ is regular. ∎
Observation 8.10.
Let $R$ be $\operatorname{F}-\emph{finite}$ and $\operatorname{F}-\emph{injective}$ with quotient singularity. Then $R$ is regular, provided $\operatorname{id}_{R}(\operatorname{Hom}_{R}({{}^{\operatorname{F}_{n}}\!}R,{{}^{\operatorname{F}_{n}}\!}R))<\infty$.
Proof.
Since $R$ is of quotient singularity, $\operatorname{Hom}_{R}({{}^{\operatorname{F}_{n}}\!}R,{{}^{\operatorname{F}_{n}}\!}R))$ is maximal Cohen-Macaulay. In view of [18, Prop 3.2], the ring is Gorenstein. But, Gorenstein and $\operatorname{F}-\emph{injective}$ are $\operatorname{F}-\emph{pure}$. Now, use the previous observation. ∎
There are some situations for which $R$ is summand of ${{}^{\operatorname{F}_{n}}\!}R$, and so the above observations can be applied for them, as well.
9. On the injective dimension of tensors
Recall that:
Question 9.1.
(See [14, Question 4.2]). Suppose $R$ is $d$-dimensional local Cohen-Macaulay of minimal multiplicity with canonical module. Let $M$ be a maximal Cohen-Macaulay module so that ${M}\otimes{M}^{\dagger}\cong\omega_{{R}}$. Then is $M\cong R$ or $M\cong\omega_{{R}}$?
The following drops the minimal multiplicity from [14, Corollary 4.5]:
Proposition 9.2.
Suppose $R$ is $1$-dimensional local Cohen-Macaulay with canonical module. Let $I$ be a reflexive ideal so that ${I}\otimes{I}^{\dagger}\cong\omega_{{R}}$. Then $I\cong R$.
Proof.
Recall that $0pt(I)>0$, as otherwise $R/\mathfrak{m}\subseteq I\subseteq R$ which is impossible, as $0pt(R)=1$. By dagger duality we mean dual with respect to $\operatorname{Hom}_{R}(-,\omega_{{R}})$. Let $x$ be $R$-regular (and so $I$-regular). Then it is easy to see
$E_{\bar{R}}(k)=\omega_{\bar{R}}=\bar{I}\otimes\bar{I}^{\dagger}.$
Take another duality gives us
$\bar{R}=\operatorname{Hom}(E_{\bar{R}}(k),E_{\bar{R}}(k))=\operatorname{Hom}(\bar{I}\otimes\bar{I}^{\dagger},\omega_{\bar{R}})=\operatorname{Hom}(\bar{I},\bar{I}^{\dagger\dagger})=\operatorname{Hom}(\bar{I},\bar{I})$
Since $I$ is reflexive, it is second syzygy and so $\bar{I}=\frac{I}{xI}\hookrightarrow\oplus\bar{R}$. In particular, $\bar{I}$ is torsionless and so
$0=\operatorname{Ext}^{1}_{\bar{R}}(\operatorname{Tr}(\bar{I},\bar{R}))=\operatorname{Ext}^{1}_{\bar{R}}\big{(}\operatorname{Tr}(\bar{I}),\operatorname{Hom}(\bar{I},\bar{I})\big{)}$
Recall that $0pt(\bar{I})=0$. Then, thanks to [22, Proposition 3.3(2)], this gives $\bar{I}$ is free over $\bar{R}$. Consequently, I{I} is free over R{R}. ∎
Proposition 9.3.
Suppose $R$ is $d$-dimensional local Cohen-Macaulay with canonical module. Let $M$ be a totally reflexive module so that ${M}\otimes{M}^{\dagger}\cong\omega_{{R}}$. Then $M\cong R$.
Proof.
By Auslander-Bridger, $0pt(M)=\dim R$. Recall that things behave well with respect to reduction by regular sequences. So, by standard reductions, the desired claim reduces to the $1$-dimensional case. Now, apply the previous result. ∎